OpenStudy (anonymous):

how do i solve (1.00+x)^2 divided by (1.00-x)^2

OpenStudy (phi):

you can't "solve" this unless you have an equation. is it set equal to something?

OpenStudy (anonymous):

its equal to 5.50

OpenStudy (phi):

you mean \[ \frac{(1+x)^2}{(1-x)^2} = 5.5 \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

I would multiply both sides by (1-x)^2 then expand each side (1+x)(1+x) = 5.5(1-x)(1-x) can you do that?